OpenStudy (benzo):

h11p1 .. q3 any help please?

OpenStudy (anonymous):

Lab 10 1/(R*C) R/L L 800000n C 20000p for last question draw a series combination of L, C and R network L - 1.6u C - 16221910p R - 10

OpenStudy (anonymous):

H11P1 a) 1/(RC*C) + RL/L b) (thats the formula = 1/(2*pi*f)^(2)*L) c) range of dis ans within 4 and 10 d) (thats the formula = 1/(2*pi*f)^(2)*L) e) range of dis ans is from 15 and 20 Note that if u doesnt get d ans in dat range ten extend or reduce the range :)